RC Meeting, Monday, October 11 , 2021

From HallCWiki
Jump to: navigation, search


RC Meeting Information

RC Daily Meetings are at 10 am, remote https://bluejeans.com/861439920

Accelerator Status

  1. For the most part we have had smooth beam delivery over the weekend.
    • There was a PSS issue that was caused by an IOC failure that took us down for about half a shift
    • After the PSS issue was resolved it looks like an administrative limit in the MCC software prevented high current in the injector (MCC will investigate to see what was the cause of the bad setting).
  2. Setup for Hall A will start around 4PM today.

Hall C Status

  1. Thanks to our shift crews we have been very productive when beam was available and we have accumulated ~9500 pion so far in this configuration.
  2. The plan is to change to the next configuration after this meeting https://logbooks.jlab.org/entry/3921453

Current Run Plan

  1. Continue data taking until shortly after 10AM:
    • Production data on LH2 target with 45 uA beam at kinematics (SHMS: 8.035 GeV/c, 5.505 deg | HMS: -1.026 GeV/c, 47.04 deg)
  2. Rotate the HMS out to achieve Q2=8.5 GeV setting
    • Determine acceptable rate in the spectrometers and maximum current
    • Start production data collection

Current Issues

  1. Replace flow switches for magnets: next opportunistic access (Joe Beaufait) ( PENDING )
  2. Water flow to the dump aperture needs to be replaced: needs 4+h and Steve Suhring believes that it can wait until the SAD ( PENDING )
  3. Target ladder camera was identified to be bad: Joe Beaufait will purchase a replacement in the next funding cycle ( PENDING )
    • Still need to follow-up with Joe about when the purchase can be done.

Special Requests

Useful Links